OpenStudy (anonymous):

16. How many solution does the following system have? 5x-2=y 3x+1=4y

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

oh wait I forgot to put the answers done

OpenStudy (anonymous):

a. one solution b. two solutions c. infinitely many solutions d. no solutions

OpenStudy (anonymous):

@wio come backkk D:

OpenStudy (anonymous):

Okay, do you know how to solve it?

OpenStudy (anonymous):

no :(

OpenStudy (danjs):

simple way to check, first put them both into ax+by = c form

OpenStudy (danjs):

if you can multiply one equation by a constant number and get the same left hand side, the lines are parallel (no solution) if you can multiply and get the same left and right sides, they are the same exact line, dependent, and have infinitly many solutions

OpenStudy (anonymous):

Substitute:\[ 5x-2=y \]Into\[ 3x+1=4y \]for \(y\), to get:\[ 3x+1=4(5x-2) \]Does this make sense?

OpenStudy (anonymous):

okay guys.. it makes sense but like I put it in on the desmo and got something but does it mean it would be a many solutions?

OpenStudy (danjs):

5x - 2 = y 3x + 1 = 4y put into ax+by = c form -------------- 5x - y = 2 3x - 4y = -1 -------------- I don't see straight away , a number you could multiply one of the equations by to get the second left side, or both sides, so it probably has ONE solution, and you need to solve for it using one of the few methods.

OpenStudy (anonymous):

the easiest was is to sketch them to see whether they interest. if they do not each other then it means there is no solution.

OpenStudy (anonymous):

*way

OpenStudy (danjs):

If you solve and get something like 10 = 20, FALSE then it is inconsistent, there are ZERO solutions, parallel lines or if you get 10 = 10 TRUE then they are dependent, the same line, infinit solutions

OpenStudy (anonymous):

*intersect

OpenStudy (anonymous):

ohhhh

OpenStudy (danjs):

well you would need a infinite sized window to make sure they didnt intersect from a graph, you need to make sure they are parallel to decide that they dont intersect, and have no solutions Parallel, Same Slope, Different Y intercepts OR The left side of ax + by is a scalar multiple of cx + dy in the second line

OpenStudy (danjs):

Example, 5x + 10y = 20 15x + 30y = 4 Parallel because 3 time the first left hand side (5x+10y) is the same as the second line left hand side, scalar multiple, parallel lines

OpenStudy (danjs):

If the entire equation is a multiple, on both sides, then Same line Infinite solutions, overlapping, dependent IE 5x + 10y = 20 15x + 30y = 60 3 times the first equation is the second equation, they are the same line, infinite solutions

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):



OpenStudy (danjs):

You can always just try to solve, and see what you get, 10 = 10 True means infinite solutions, same line 8 = 6 False means no solutions, parallel lines x=#, y=# means that is the one intersection, one solution

OpenStudy (anonymous):

ohh okay I see now

OpenStudy (danjs):

How can 2 Lines have 2 solutions? that is choice B

OpenStudy (danjs):

without bending one of them

OpenStudy (anonymous):

ohh oki thanks guys <3

OpenStudy (anonymous):

*Hint* When you only need to know the number of solutions, simply look at the problem. A first degree equation will have ONE solution: \[x+5=15\]A second degree equation will have TWO solutions. \[x^2 + 4=20\]A third degree equation will have.. THREE :) And so forth.